### v1

# Inherent and emergent liability issues in LLM-based agentic systems: a principal-agent perspective

Abstract
Agentic systems powered by large language models (LLMs) are becoming progressively more complex and capable. Their increasing agency and expanding deployment settings attract growing attention over effective governance policies, monitoring and control protocols. Based on emerging landscapes of the agentic market, we analyze the potential liability issues stemming from delegated use of LLM agents and their extended systems from a principal-agent perspective. Our analysis complements existing risk-based studies on artificial agency and covers the spectrum of important aspects of the principal-agent relationship and their potential consequences at deployment. Furthermore, we motivate method developments for technical governance along the directions of interpretability and behavior evaluations, reward and conflict management, and the mitigation of misalignment and misconduct through principled engineering of detection and fail-safe mechanisms. By illustrating the outstanding issues in AI liability for LLM-based agentic systems, we aim to inform the system design, auditing and monitoring approaches to enhancing transparency and accountability.
1 Introduction
AI agents are computer software systems capable of creating context-specific plans in non-deterministic environments (Chan et al., 2023; Krishnan, 2025). AI agents based on LLMs (aka. LLM agents, see Appendix A) exist on a spectrum of autonomy, ranging from simple tool-calling agents to generalist agents capable of planning, sourcing, critiquing, and executing their own workflow (Li, 2025). They primarily adopt an architecture with explicitly defined functioning components [^1] Also called cognitive architecture (Kotseruba & Tsotsos, 2020) at times, but the architecture alone doesn’t guarantee cognition or sense of agency. (Sumers et al., 2023). LLM-based multiagent systems (MASs) allow agents to interact, collaborate, or compete within shared environments (Fig. 1a). They are designed by combining agents with specialized roles through LLM role-playing (Shanahan et al., 2023; Chen et al., 2024a) or through integration on a software platform. Each agent handles specific subtasks based on its expertise and allocated resources (tools, data, compute, etc). MAS can be tailored to a wide range of scales and domains, from few-agent systems that simulate team decision-making in medicine (Tang et al., 2024; Kim et al., 2024) and finance (Xiao et al., 2025), to many-agent systems that mimic the population-level socioeconomic dynamics (Park et al., 2024a; Piao et al., 2025). A plausible LLM-based MAS (Fig. 1a) can possess multiple teams of interacting agents coordinated by an orchestrator (Wang et al., 2025) and behaviorally regulated by other platform agents or through a set of engineered constraints (e.g. norms) (Criado et al., 2011; Hadfield-Menell et al., 2019). While the flexibility of LLM-based MAS provides adaptability to various applications and additional affordances but also emergent risks not present in single agents (Hammond et al., 2025; Pan et al., 2025).
Figure 1: (a) A plausible LLM-based MAS deployed on an agent platform, where delegation goes from the principal to an orchestrator (agent) to different functioning agent teams. The platform also contains supporting agents. Black circles indicate teams. (b) Landscape of the LLM agentic market. LLM agents are supplied through agent-native services or agentic SaaS (agents on top of existing software apps). Agent suppliers are supported by model suppliers. Colors in (a)-(b) distinguish between agents of different types or suppliers.
Understanding the current landscape of the agentic market [^2] https://aiagentslist.com/ai-agents-map (Fig. 1b and Appendices A-B) is essential for analyzing liability. The present work uses the following terms to refer to its key components: Software platforms offering agent-native services include generalist agents that focuses on general autonomous use of computers, specialist agents which target labor intensive sectors and provide verticalized services for domain-specific workflow automation (Bousetouane, 2025), and character-infused, “hireable” agentic employees. Specialist agents are also offered directly by established software platforms as agentic SaaS to streamline their existing app services. Both agent-native services and agentic SaaS largely source models externally from model suppliers or derive their own models from open-source projects. Separately, prototypes of integration platforms and integration protocols facilitating the interaction of agents from different frameworks with third-party resources during deployment are also appearing.
Although the potential liability issues in the rapidly expanding market are prevalent, the governance of AI agents is still a nascent topic. Existing efforts are built along two streams: one focuses on establishing and refining the taxonomies of risks and harms using empirical evidence (Chan et al., 2023; 2024; He et al., 2024; Hammond et al., 2025) or differentiating the governance of agentic systems from traditional machine learning (ML) models (Cohen et al., 2024; Kolt, 2025); another focuses on understanding the interactions between humans and AI agents to build constructive principles (Zheng et al., 2023). Research in both streams used principal-agent theory (PAT) (Eisenhardt, 1989; Laffont & Martimort, 2002) as a starting point, but lacked systematic examination of how existing legal liability frameworks can effectively address various principal-agent relationships in AI systems. Moreover, AI systems are typically embedded in a sociotechincal system (Weidinger et al., 2023) such that the risks are materialized only through interacting with the environment (see example and analysis in Appendix C).
Despite the long history of PAT-based legal frameworks (Munday, 2022), their use in AI systems is still in early days. This work presents an initial attempt in extending principal-agent analysis to current LLM-based agentic systems. Our major contributions include: (i) extending the previous work on PAT for AI governance to accommodate the characteristic behaviors of LLM agents; (ii) discussion of the potential liability issues in light of the emerging market for LLM-based agentic systems; and (iii) outlining of the important directions of policy-driven method developments that support the technical governance of LLM agents.
2 Related works
AI liability
Legal scholars have treated AI as products and the harm they caused under product liability (Abraham & Rabin, 2019; Buiten, 2024; Sharkey, 2024). In the US, product liability falls under: (i) design defect, (ii) manufacturing defect, and (iii) a warning defect (inadequate instructions) (Abraham, 2012). Anyone harmed by a product can bring a claim. Product harm carries strict liability — meaning courts do not consider the care level — and affects anyone along the supply chain.
Service-caused harm usually trigger a negligence inquiry: was “reasonable care” used? If not, the service provider is liable (primary liability) and their principal can also be liable under multiple legal theories. Principals face two primary liabilities: “negligent hiring” for hiring an agent without a reasonable due diligence on its modus operandi; and “negligent supervision” for failing to reasonably monitor or control its agents. Principals face secondary liability such as vicarious liability (Sykes, 1984; Diamantis, 2023) because of their relationship (when its agents acted within the boundaries of employment). A service relationship can also fall under a contractor relationship, which does not trigger vicarious liability. Courts look at the contract and control exercised to categorize the relationship. Below, we assume the principal exercise enough control to trigger vicarious liability.
Software exhibits both “product” (movable goods) and “service” (akin to professional offering) characteristics (Gemignani, 1980; Popp, 2011). As AI agents become more autonomous, they move closer to services, their actions are more accountable due to increasing agenticness (Chopra & White, 2011; Chan et al., 2023). A negligence rule with potential vicarious liability may be more suited (Turner, 2018) for those relationships. In current legal use, this framework straddles product liability falling on the manufacturer and vicarious liability involving principals. Besides the service-product divide, law and economics (L&E) has also advocated other approaches, including risk-based liability (Geistfeld et al., 2022), fault-based liability (Buiten et al., 2023), explanation-based liability (Padovan et al., 2023), etc.
Principal-agent problems in AI/ML
Prior works that invoked the principal-agent framework in AI/ML primarily focused on the decision-making aspects in human-AI collaboration and AI safety. Lubars & Tan (2019) discussed the relation between task delegation and the principal’s preference. Hadfield-Menell & Hadfield (2019) mapped AI alignment onto the principal-agent problem and discussed the alignment issues in the incomplete contracts theory. Athey et al. (2020) considered different scenarios in allocating decision authority when the human principal and the AI agent have different capabilities. Critch & Krueger (2020) and Hendrycks (2023) considered the potential dangers of complete task delegation to agentic systems. Besides, principal-agent problems have also been considered in game-theoretic machine learning (Gan et al., 2024) and in reinforcement learning settings with two interacting agents (Ivanov et al., 2024).
Agent-oriented software systems
Agentic systems have long been proposed as a canonical approach for software design (Jennings, 2001; Zambonelli et al., 2003). Their core advantages are the scalability, flexibility, and the ability to perform complex tasks through task decomposition. LLM agents (Li, 2025) can be configured by text instruction and can interact with external resources to achieve enhanced capabilities than LLMs in reasoning, tool use, memory, planning, and can be personalized according to the role-playing capability of the base model (Shanahan et al., 2023; Chen et al., 2024a). LLM-based MASs can use verbal communication protocols to facilitate collaboration and engage in debates (Tran et al., 2025), and the protocol topology is a key for their efficient scaling and behavioral control (Qian et al., 2025).
3 Principal agency and liability
PAT examines the relationship where one party, the principal, delegates authority to another, the agent, creating three fundamental challenges (Eisenhardt, 1989; Laffont & Martimort, 2002):
  1. Adverse selection (aka. hidden information problem) occurs when agents possess more information than principals about their abilities or efforts. Adverse selection is a type of information asymmetry also known as the hidden information problem (Akerlof, 1970).
  2. Moral hazard (aka. hidden action problem) occurs when agents take greater risks than principals would prefer because they do not bear the full consequences (Arrow, 1963). Moral hazard is also a type of information asymmetry.
  3. Misaligned interest (aka. conflict of interest) can manifest between principals and agents in four ways (Jensen & Meckling, 1976): principal-agent collusion against third parties, principal-third party collusion against agents, agent-third party collusion against principals, or agents simply pursuing self-interest independently. Misaligned interests lead to agency cost, which is associated with information sharing, monitoring of the agent, etc (Fama & Jensen, 1983).
Each of these problems has legal and economic solutions. Legal solutions include fiduciary duties (duty of care, duty of loyalty, etc.). Those duties provide principals with a legal recourse against agents who breached those duties. However, legal enforcement is probabilistic and slow (Levmore, 1990), so principals often prefer to use economic mechanisms. Economic solutions depend on the problem. Principals can address some hidden information problems by creating separating equilibria that force agents to reveal information about themselves, that is, “signals”; signals include credentials, warranties, or performance histories that distinguish high-quality from low-quality agents (Spence, 1973). Principals mitigate a hidden action problem through monitoring (i.e. direct observation of agent behavior) and then linking compensation to observable effort, or bonding arrangements in which the principals tie remuneration to outcomes (bonuses), thereby aligning financial incentives (Holmström, 1979). A principal can address conflict of interest by realigning incentives through carefully designed contracts (deferred compensation), creating organizational structures that promote incentive realignment (profit sharing), or leveraging reputation mechanisms. A principal has incentives to monitor their agents. Primary and secondary liability exposures provide additional incentives to ensure that it oversees that its agents complete their tasks using reasonable care.
4 Inherent liability issues in single agents
Contemporary approaches to the governance of AI agents (Kampik et al., 2022; Chan et al., 2023; 2024; Kolt, 2025) resort to PAT, where the principal, the human or company, delegates a task or goal to the AI agent, based on a mutual agreement. Yet LLM agents still cannot satisfy all criteria of a normal agent (Perrier & Bennett, 2025) in PAT, therefore, the agency gap can create additional liability issues or liability gaps (Turner, 2018).
4.1 Artificial agency
PAT requires a clarification of the agency relationship, which remains a hotly debated interdisciplinary topic for LLM-based systems (Shavit et al., 2023; Dai, 2024; Barandiaran & Almendros, 2024; Dung, 2024; Perrier & Bennett, 2025; Mattingly & Cibralic, 2025; Butlin, 2025; Das, 2025). AI researchers often take an operational view of artificial agency, such that it is possible to quantify and compare the mental state characteristics of and between AI systems through external interrogation (Baird & Maruping, 2021; Chan et al., 2023; Miehling et al., 2025). Representative caveats of artificial agency for LLMs and LLM agents include:
  1. Instability: Behavior varies with the same or paraphrased but meaning-consistent prompt on different trials (Loya et al., 2023).
  2. Inconsistency: Behavior is sensitive to distracting contextual information or affected by sentiment and adversarially designed prompts (Jain et al., 2023; Maus et al., 2023; Zhuo et al., 2024).
  3. Ephemerality: The complexity of behavioral sequence is restricted by the context window length because of the lack of effective memory mechanisms (Maharana et al., 2024).
  4. Planning-limitedness: Construction of executable plans hinges on accessible environmental feedback which is task-limited (Kambhampati et al., 2024; Wang et al., 2024b; Chen et al., 2024b).
These caveats finds similar manifestations in behavioral economics (Thaler, 2016).
Liability from flawed agency
Because an agency relationship requires both principals and agents agree to the relationship, current LLM agents cannot yet form an authentic relationship of such because of their flawed agency (Barandiaran & Almendros, 2024; Perrier & Bennett, 2025). Voluntary relationships usually indicates that both parties benefit. However, because rationality is generally not a built-in goal in developing LLM agents (Macmillan-Scott & Musolesi, 2024), their lack of consent (from agents) means that agent suppliers may face liability or risks they have not considered. That is why scholars have consistently argued that agent failures should be treated as product liability, and the allocation should reflect that (Abraham & Rabin, 2019; Buiten, 2024; Sharkey, 2024). In such a situation, the Coase Theorem suggests that the allocation of liability and risks is best left to contracts but before the parties can do that, legislators and courts must clarify rights and liabilities (Coase, 1960).
4.2 Task specification and delegation
Principals delegate tasks to their agent usually either because the principals lack the resources or the expertise to deal with the tasks (Castelfranchi & Falcone, 1998). For LLM agents, the principal provides as input to the LLM a task specification that contains instructions on the nature and procedure of task execution, available resources, potential ways to overcome hurdles, the principal’s preferences, etc. Such a procedure is subject to
  1. Task underspecification is often present because it is impossibly costly to fully anticipate all possible scenarios to put into the specification. Moreover, the same agent may be appropriate for one task but not another; therefore, task specification — the equivalent of job design in organizational theory (Oldham & Fried, 2016) — should balance the tradeoff between agent capability and task complexity (Hadfield-Menell & Hadfield, 2019). This incompleteness leaves the door open to undesirable outcomes such as negative side effects.
  2. Risky delegation The principal tends to forgo the delegation of very high-stakes tasks because of their severe adverse consequences and the unreliability of agent behaviors (Lubars & Tan, 2019). The risk of delegation can be reduced by the amount of repeated feedback the principal provides to ensure alignment (Jiang et al., 2024).
Liability from task misdelegation
When selecting the agent (and the tasks) to delegate, principals are expected to carry out their due diligence. If they fail to take reasonable precautions when selecting agents, principals retain liability for negligent hiring (Camacho, 1993) and negligent retention. At the best of time, delegation of tasks raises information concerns because the principal may not have authorized the agent to share information. In the LLM context, the sharing of information raises concerns of copyright and trademark (e.g. duplication of protected documents), trade secret (sharing information outside the system), privacy (e.g. transferring data governed by the General Data Protection Regulation or California Consumer Privacy Act), etc. That is why many tasks are not delegable. A human agent may be able to dissociate the difference between delegable and non-delegable tasks based on the sensitivity of the information, while an AI agent may not without clear instructions (Hadfield-Menell & Hadfield, 2019).
Figure 2: Examples of interaction patterns between the principals and single agents or multiagent systems (MASs). Each MAS in the coordination or collaboration (collab.)/competition (compet.) pattern is enclosed within a box.
4.3 Principal oversight
Human oversight is a costly endeavor for many AI applications yet it remains the gold standard in existing AI governance principles (Sterz et al., 2024; Cihon, 2024). Principal-agent problems suffer from information asymmetries, which are usually resolved through monitor and incentive realignment via reward or punishment. However, monitoring an AI agent requires designs that allow the principal to observe and understand what the agent is doing. Principals who prefer to realign the AI agent’s incentives (Everitt et al., 2021) with their own would have to understand what “motivates” AI systems or assume that AI responds to human-style incentives (Ratliff et al., 2019). At the moment, quantifying misalignment remains challenging, especially for capable and general AI systems (Anwar et al., 2024). Moreover, the quality of principal oversight can be threatened by a spectrum of behaviors LLMs inherit from their training, such as:
  1. Sycophancy refers to the tendency of AI systems to provide responses that the evaluator would prefer in favor of improving the answer (Perez et al., 2023; Sharma et al., 2024), exploiting the evaluator’s cognitive biases (e.g. susceptibility to flattery) rather than correctly performing their duty.
  2. Manipulation refers to the ability of LLMs to influence their principals (Campedelli et al., 2024; Burtell & Woodside, 2023; Carroll et al., 2023), towards ends that are non-welfare maximizing to their principals.
  3. Deception refers to the tendency of AI systems to induce false beliefs (Park et al., 2024b; Scheurer et al., 2024; Lang et al., 2024), reinforcing the information asymmetries between the AI agent and its principal.
  4. Scheming refers to the strategic behavior of AI systems to harbor alternative and potentially harmful motives from alignment with the principal during post-training, leading to fake alignment (Greenblatt et al., 2024; Balesni et al., 2024)
These behavioral patterns also manifest in real-world scenarios (Blonz, 2023). All of them can upend the principal-agent asymmetry such that the principal will not be able to reliably monitor the agent’s behavior or provide informative feedback.
Liability from compromised oversight
Principals that fail to exercise oversight over agents that committed unlawful acts are subject to primary liability under a negligent supervision theory or negligent retention (Cavico et al., 2016) depending on the nature of the act. The principal faces liability because if the principal had exercised reasonable investigation, it would have known that the acts (or omissions) of its agents would harm third parties or create unreasonable risks.
5 Emerging liability issues in MASs
MASs have flexible designs (Fig. 2), so agency relationships and the associated principal-agent problem can occur at different levels, similar to the functioning of a firm (Fama, 1980). The challenges in governing single LLM agents are compounded with multiple participating agents and complicated by emergent issues with no equivalent in the single-agent setting. An orchestrator (agent) in an MAS functions as a local manager, and directs the work execution of a number of agent workers/teams with different expertise but imposing liability only on the orchestrator does not incentivize the improvement of the subagents. Liability plays an important role when market mechanisms such as reputation do not function well (e.g. lack of repeated interactions, monopoly power in an agentic space, etc). For compatibility with PAT, we consider here only the MASs where the agent roles/tasks are assigned before at start. We discuss three aspects in the following subsections.
5.1 Role and agency allocation
Constructing an LLM-based MAS involve role (or task) allocation (Campbell & Wu, 2011; Guo et al., 2024; Tran et al., 2025), which implicates allocated agency, where the single agents can act on their own to accomplish the goals defined by assigned task and resources. Because of the role-playing capability (Shanahan et al., 2023; Chen et al., 2024a), current LLM-based MASs generally adopts a role-centric approach, executing role allocation alongside its associated task (Guo et al., 2024; Tran et al., 2025). This approach provides an interpretable division of labor and can directly mimic interactions in teams of humans. Alternatively, role allocation can be self-organized such as in the deployment of subagents, which also need not be stationary. Role allocation is affected by the nature of LLMs’ flawed agency (Section 4.1), leading to potential downstream issues:
  1. Influenceability: Agency of individual agents in an MAS can be enhanced or reduced through communication with other agents in a cooperative MAS (He et al., 2025), therefore triggering agency shift or unanticipated conducts.
  2. Distributedness: The distribution of agency to different agents in an MAS leads to task specialization and latency which can trade off against performance and speed (Mieczkowski et al., 2025).
  3. Diminished control: In a hierarchical MAS, subagents are more separated from the principal than the head agent, therefore may be harder to directly control or monitor. The principal is more prone to manipulation from the head agent.
Liability from agent misallocation
These concerns make task allocation the principal’s most important decision. First, the principal must decide which tasks can be delegated. As the number of allocations increases, the agency cost also increases: each new allocation must offer marginal benefits that justify its marginal costs. L&E favors assigning liability on the entity that can avoid the accident at the lowest cost (Calabresi & Melamed, 1972; Carbonara et al., 2016). Second, MASs expose principals to the risks their (sub)agents take. The principal must understand the risks associated with each task and with each agent. Then, it must assess which tasks can be delegated without surpassing its risk tolerance and to whom. However, the principal faces information asymmetries: it might speculate about the expected cost-benefit of each (sub)agent. It faces liability for not being careful assigning a task to an agent but usually not for allocating too many tasks (Carbonara et al., 2016). Finally, if the system is so complex or opaque, courts may decide that the harm “speaks for itself”, inferring carelessness from the harm — turning a negligence rule into a strict liability rule (Fraser et al., 2022; Casey, 2019). Explainability can therefore become even more necessary.
5.2 Operational uncertainty
As LLM-based MASs become more complex, oversight becomes increasingly challenging. A human overseer may only handle direct communications with the head agent while the subsequent interactions between the head agent and the subagents are initiated autonomously among themselves. Although the organizational hierarchy and communication protocol can benefit from minimal human involvement such as in the coordination structure in Fig. 2, multiagent interactions via LLMs create additional challenges depending on the agent architecture and task (Pan et al., 2025).
  1. Failure cascade refers to the scenario where the downstream agents can have increased vulnerability than the upstream agents in a MAS with coordination structure (Peigne-Lefebvre et al., 2025).
  2. Rogue agent refers to the event that an agent in confusion introduces communication noise which can lead to system failure (Barbi et al., 2025).
  3. Agent collusion refers to the collaboration between agents that negatively impact others (Fish et al., 2024; Lin et al., 2024).
A promising direction for minimizing operational uncertainty from misbehaving agents is to instigate corrective mechanisms and foster a norm-based governance (Hadfield-Menell & Hadfield, 2019; Kampik et al., 2022), where the norm is defined through spontaneous and engineered social interactions between agents (Trivedi et al., 2024).
Liability from operational uncertainty
When multiple LLM agents interact autonomously, the attribution of responsibility becomes blurred as decisions emerge from collective behaviors rather than individual actions. A software provider usually bears the liability for the harm their LLM agents cause. In heterogeneous MAS, courts may apply separate liability or distribute liability among suppliers according to the harm contribution. Courts do not favor this approach because it is complex to estimate, so the parties would benefit from contractual clarification between all the (sub)agents (e.g. contract liability based on value). Instead, when causes cannot be disentangled and assigned, courts revert to join and several liability — meaning that each party is liable for the full harm and can be sued individually, which expands liability to the (sub)agents — a tempting approach to avoid complex litigations and battle of experts when MAS are involved (Custers et al., 2025). Courts may decide that MAS-related behaviors such as failure cascade, rogue agent, or agent collusion tendencies lead to third-party harm too often. So, courts may elevate the use of MAS to a “risky activity” and use a strict liability rule (Čerka et al., 2015).
5.3 Platform integration
Agent behaviors and interaction patterns are constantly changing. Emerging LLM-based MASs will likely feature supplier-dependent agent frameworks, which will follow somewhat different safety protocols (Fig. 2). At the moment, efforts to integrate different agent frameworks analogous to traditional software integration (Bass et al., 2021) are still lagging behind but are expected to ramp up due to the market growth. The motivation to integrate LLM agents is to enhance system capability by unifying disparate supplier frameworks, which may include privileged access to customized agentic components (e.g. unique databases, fast memory, etc). Overall, integration serves the needs of the user (i.e. principal) by balancing the advantages of different agent frameworks. At the moment, the potential benefits of integration platforms for LLM agents include:
  1. Platform oversight refers to measures on platform to provide users with enhanced multiagent security through a security-guard agent (Xiang et al., 2024), collusion mitigation mechanism (Foxabbott et al., 2023), detection and suppression of copyright infringement (Liu et al., 2025a) or privacy leakage. These oversight mechanisms are not necessarily baked into the design of individual agents but are provided through the integration platform.
  2. Platform teaming refers to the formation of agent teams on an integration platform through user-defined or ad hoc protocols (Mirsky et al., 2022; Wang et al., 2024a) that enhances cooperation among homogeneous or heterogeneous agents.
Liability from mismanaged platforms
In an integrated multiagent platform, the platform could carry some liability depending on the level of control they exercise (Gabison & Buiten, 2020; Lefouili & Madio, 2022). For example, control might encompass the behavioral monitoring of the individual agents operating on the platform. Because the platform intrinsically involves multiple principals and multiple agents, the principals may be liable for engaging in collusive behavior but, in rare occasions, platforms have faced liability for incentivizing others (e.g. copyright infringement [^3] MGM Studios, Inc. v. Grokster, Ltd., 545 U.S. 913 (2005), intentional interference444hiQ Labs, Inc. v. LinkedIn Corp., 938 F. 3d 985 (9th Cir. 2019)).
6 Policy-driven technical development
Improving the transparency and accountability of LLM-based agentic systems that supports in-depth failure analysis and liability attribution requires concomitant developments in agent behavior and system management, as discussed below:
Interpretability and behavior evaluations
Tracing an AI agent’s actions (Lu et al., 2024) can be the basis for establishing whether it took reasonable care (Price et al., 2019; Choi, 2020) and therefore the evidence for liability claims. Prior works on the interpretability and faithfulness in LLM reasoning Lyu et al. (2023); Wei Jie et al. (2024) and dialogue generation (Tuan et al., 2021) may eventually also be used to assist in the analysis of agent behavior. More generally, these kinds of evaluations may help understand unreliable behavioral patterns and help diagnose and refine the design bottlenecks in LLM agents. To better ground notions of reasonable care for AI agents, method developments should prioritize decomposing complex multiagent interactions into interpretable causal mechanisms, leveraging causal abstraction frameworks Geiger et al. (2024) to create faithful, human-intelligible representations of agent interactions that preserve essential causal relationships while abstracting away unnecessary details. Additionally, formal verification approaches (Zhang et al., 2024) may be able to detect and prevent potential failure modes in agent interaction and decision-making.
Reward and conflict management
Some application settings of LLM-based MASs aim to mimic the functioning of human teams and organization (Xie et al., 2024). Agentic systems can learn from existing organizational theory (Mitnick, 1992; Vardi & Weitz, 2016) to improve the design and architecture given the flawed agency of its components. The most relevant aspects include systems for managing reward and conflict between agents. It has been shown that “verbal tipping” (Salinas & Morstatter, 2024) can provide incentive in instruction to improve LLM performance. Moreover, managing knowledge conflict in LLMs (Xu et al., 2024) has been the most related area. To manage generic conflicts between LLM agents, a credit system (Thomas et al., 2017) of refusal and sanction based on agent IDs may be beneficial (Chan et al., 2025). This could include adaptive trust scoring from evaluation of domain-specific expertise, and an arbitration protocol to adjudicate conflicts such that the arbiter holds the right to refuse action from a frequently misbehaving agent. While an agent scoring system would need to be supported across the relevant AI agent integration platforms (Fig. 1a), it would be essential to balance individual agent utility and cooperation within the system.
Misaligment and misconduct avoidance
As discussed in Section 4.3, LLMs and the agents based on them may act in misaligned ways, with behaviors ranging from sycophancy and deception to scheming. While AI model suppliers may be able to partially reduce these problems through effective detection and behavior steering, which has been demonstrated on LLMs (Rimsky et al., 2024; Goldowsky-Dill et al., 2025; Williams et al., 2025). In LLM-based agentic systems, the equivalent tasks could be carried out using separate LLM agents through observing and analyzing other agents’ behavior, such as using the theory-of-mind capability (Street, 2024). The MAS can include agents with a specially finetuned base model (Binz & Schulz, 2023) as a warden (agent) for deception mitigation. Similar approaches can also be used to suppress other agent misconducts such as generation of harmful or copyright-protected content using the warden to filter through key tokens and phrases that can induce such behavior. The adaptive approach carried out by a separate agent can compensate for the limitations in existing model-level approaches such as machine unlearning (Bourtoule et al., 2021), which is suffering from limited effectiveness and tradeoffs with other model capabilities (Cooper et al., 2024; Liu et al., 2025b).
7 Conclusion
We examined liability issues arising from LLM-based agentic systems thorough analyzing distinct aspects of the agentic AI ecosystem according to the principal-agent theory. Despite that a varieties of issues are yet to present themselves in the real world, the growing evidence and demonstrations in simulated scenarios can inform their potential impact at the societal scale, which we built on in our prospective study. Our work shows that besides increasing agency, disruption in other aspects of the principal-agent relationship can also lead to liability incidents. Ultimately, the materialization of liability issues in reality will be dominated by the more frequent use cases of agentic systems in each industry sector. Our analysis enriches existing contextualization of AI risk (Chan et al., 2023; Hammond et al., 2025) and demonstrates the explanatory power of the behavior-centric approach to translating frontier AI research into tangible knowledge to inform legal analysis and policy.
Ethics Statement
The design and deployment of LLM-based agentic systems raise significant ethical and legal considerations regarding liability attribution. This work acknowledges the complex interplay between supplier responsibility, user actions, and the emergent behaviors of semi-autonomous systems powered by LLMs. We recognize that traditional liability frameworks may inadequately address scenarios where AI agents make consequential decisions with incomplete human oversight. Our analysis aims to contribute to the discussion on appropriate liability for all sides related to LLM-based agentic systems that balances current evidence from system behavior and the available legal framework. We have considered liability in different deployment environment with the aim of informing policy discussions and technical developments to make liability more traceable in AI systems that are complex and extensible, but prone to misbehavior and failure and lack in self-control.
Acknowledgments
We thank the organizers at the AI Governance at the Crossroads symposium held at Berkeley, California in February, 2025. We thank Micah Carroll at UC Berkeley for information about LLM behaviors and AI safety and helpful comments on the manuscript.
Appendix A Definitions and their scopes
We clarify here the definitions and scopes for some of the key terms used throughout the text. They are not meant for a complete characterization of these terms but are primarily aimed at illustrating their relationships in the context of this work.
  1. An AI agent is a generic term referring to a software agent powered by any form of artificial intelligence (Krishnan, 2025). An equivalent definition is provided in Section 1.
  2. An LLM agent refers specifically to an AI agent powered by at least one LLM as the central component that executes planning, initiates and coordinates the agent’s actions, etc.
  3. An AI system is an umbrella term defined in the EU AI Act [^5] https://artificialintelligenceact.eu/article/3/. When deployed, an AI system can operate with various levels of autonomy and can create recommendations and content, make predictions and decisions that influence the environment. The system in the term indicates that the AI is not acting in isolation, but assisted by the surrounding infrastructure such as cloud computing, databases, and user interfaces, which are integral to the AI system and essential for its use. An AI system can be agentic or non-agentic.
  4. An agentic system, also known as an agenic AI system Shavit et al. (2023), is a type of AI system that contains a level of agency such that it can carry out actions on its own in pursuit of a goal. An agentic system can include a single agent or multiple agents acting in coordination, competition, cooperation (or collaboration). An AI agent is a component of an agentic system.
  5. Analogous to the previous entry, agentic market is the segment of AI market represented by the suppliers and buyers of agentic AI systems.
  6. An agent platform provides resources and toolkits to construct, configure as well as deploying AI agents (Fig. 1a). AI agents from distinct suppliers can operate on a (software) integration platform, where they can interact with each other, with third-party data sources, proprietary APIs, etc. AI agents on any deployment platform may subject to compliance governance and can receive protection against cyberthreats or malfunction from the software infrastructure there.
Appendix B Types and examples of existing LLM agent suppliers
We present here preliminary examples for elements of the agentic market that are currently available.
B.1 Agentic software as a service (SaaS)
B.2 Agent-native service
Suppliers of generalist agents
  1. OpenAI Operator: operator.chatgpt.com
  2. Google DeepMind Project Astra: deepmind.google/technologies/project-astra
  3. Please/MultiOn: please.ai
  4. Manus: manus.im
Suppliers of specialist agents
  1. Sesame (www.sesame.com) offers voice AI agents for different domains.
  2. Contextual (contextual.ai) offers specialized AI agents with advanced retrieval-augmented features.
  3. Devin (devin.ai) offers AI agents for coding.
  4. Sierra (sierra.ai) offers AI agents tailored for different types of customer services.
  5. Health Force (www.healthforce.ai) offers human resources AI agents to handle digital text processing tasks in healthcare systems.
  6. Zenity (www.zenity.io/) offers security-focused AI agents.
Suppliers of character-infused agents
  1. Artisan: www.artisan.co
  2. Sintra: sintra.ai
B.3 Elements of agent integration platforms
  1. Anthropic Model Context Protocol: modelcontextprotocol.io
  2. IBM Bee AI: beeai.dev
Appendix C Principal-agent analysis of an LLM-related legal case
Refer to caption In 2023, a US Federal Judge reprimanded two lawyers and their law firm for acting in bad faith and making misleading statements to the court. Their crime? They trusted ChatGPT. ChatGPT made up cases, the lawyers failed to notice, but the judge did. The judge ordered the lawyers to produce the cases. The lawyers “doubled down and did not begin to dribble out the truth” for another few weeks. The court punished the lawyers for the actions of ChatGPT and their law firm in the process. ChatGPT (and its provider OpenAI) did not bear any responsibility because the lawyers were the ones who had a duty to provide accurate information to the court. The lawyers responded that: “We made a good faith mistake in failing to believe that a piece of technology could be making up cases out of whole cloth.” These lawyers and their principal learned not to trust AIs.
Figure 3: Principal-agent analysis of Mata vs. Avianca, Inc.
The case of Mata v. Avianca, Inc. [^6] Mata v. Avianca, Inc., 678 F. Supp. 3d 443 (SDNY 2023) (above) was much discussed in the public media in 2023 as an early incident involving the use of LLMs. The delegational structure of major entities involved in the case is (Fig. 3)
[…]
The case features a few different principal-agent relations, starting from Roberto Mata, the plaintiff. The segment relevant for current discussion concerns the law firm that hired the two lawyers representing Mata. A part of the tasks that the lawyers do was delegated to the LLM, ChatGPT, which is the subagent of the law firm. The law firm can control the tools that the lawyers use through the employment contract, but ChatGPT was not excluded explicitly. The lawyers defended the fake cases generated by ChatGPT and therefore was ruled by the judge to be in bad faith. The law firm also received punishment alongside its employees.
The prompting method that the lawyers used was not explicitly discussed in the court proceeding. ChatGPT is treated as a subagent that carries out the lawyers’ task request. It was largely regarded as an instance of LLM hallucination (or confabulation) in the news media rather than deception. The judicial verdict was finalized based on the behavior of the lawyers. The legal case highlights the shortcomings of current legal frameworks and the technical gap to fully comprehend behaviors of AI systems in the presence of their environment (involving humans) and make use of these evidence in attributing blame.